Use the given data to find the minimum sample size required to estimate a population proportion or percentage.

Margin of​ error:

0.080.08​;

confidence level

9090​%;

ModifyingAbove p with caretp

and

ModifyingAbove q with caretq

unknown
Answer #1

solution:-

given that margin of error E = 0.08

the value of 90% confidence from z table is 1.645

here we have not p and q values so

take p = 0.5 the q = 1- p = 1 - 0.5 = 0.5

formula for sample size

=> n = (z/E)^2 * P * Q

=> n = (1.645/0.08)^2 * 0.5 * 0.5

=>n = 106